I am confused on how to do elimination. I know how to the basic kind like 5x-2y=4, and x=2y=8. But how do you do it when the signs don't match up like 7x-4y=-10 and 4y=x-2. I am so confused on that part of elimination. Also how do you do x=2y=-2 and 4x=2y=-17. Here's what I have done for the first one.
7x-4y=-10
x=4y=2
8x=-8
x=-1 and y=3/4
Please help me!

Answer by rapaljer(4671)

You have a problem in the way you are stating the problems. In many of your examples, you are giving two equal signs in a given equation. This cannot be! We can't tell what the problem is supposed to be. However, to answer your question anyway, if the equations are NOT in standard form Ax + By = C, where the x and y terms line up nice and neat, IF they don't line up, then you should re-arrange the equation so that it DOES line up nicely!

For example,
5x-2y = 4
x = 2y - 4

In the second equation things don't line up, so subtract 2y from both sides to make it line up in standard form. Rewrite the first equation as is, and rearrange the second equation so it looks like this:
5x-2y = 4
x-2y = -4

Then you can solve it accordingly. By the way, if one of the equations is in the form x = 2y -4, you might want to try the substitution method.

R^2 at SCC